OpenStudy (anonymous):

What is the area or an equilateral triangle with side of 10 inches and height of 7inches? A. 44sq. In. B. 20sq.in. C. 70sq.in. D. 35 sq. in

OpenStudy (anonymous):

\[A _{\Delta} = \frac{ bh }{ 2 }\] Multiply the side-value to the height value, then divide all that by 2 to get your answer.

OpenStudy (anaise):

hi

OpenStudy (crabbyoldgamer):

B*h/2 = 10*7/2 = 35 square inches